OpenStudy (anonymous):

Which trinomial is a perfect square trinomial? (Points : 1) y^2 + 25y + 200 y^2 + 9y + 81 y^2 + 14y + 49 y^2 + 6y + 36

OpenStudy (jadeishere):

cancel out the first one and the second and the last. So, y^2 + 14y + 49 is a perfect square trinomial, if I remember correctly

OpenStudy (whpalmer4):

A perfect square trinomial will have the following form: \[(x+a)(x+a) = x^2 + 2ax + a^2\] In other words, the coefficient of the \(x\) term, divided by 2 and squared, will give you the constant term. For example, the first one is not a perfect square, because \((\frac{25}{2})^2 \ne 200\) You can also use the discriminant of a quadratic in the form \(ax^2 + bx + c = 0\) to assess whether you have a perfect square. If \(\Delta = b^2-4ac = 0\), you have a perfect square trinomial. Again, using the first one: \(a = 1, b = 25, c = 200\) \[\Delta = b^2-4ac = (25)^2 - 4(1)(200) \ne 0\]so it is not a perfect square.